OpenStudy (anonymous):

Find the zeros of each function. State the multiplicity of multiple zeros. 1. y=(x-2)(x+4) 2. y=(x+1)(x-8)(x-9) 3. y=x(x+1)(x+5)

OpenStudy (anonymous):

first one set \[x-2=0\]solve for \(x\) in your head then set \[x+4=0\] and again solve for \(x\) others are the same procedure

OpenStudy (anonymous):

wait I don't understand.. I'm not even sure what the question is asking

OpenStudy (anonymous):

the question is asking which values of \(x\) would you plug in to get zero

OpenStudy (anonymous):

in other word, it is asking to solve \[(x-2)(x+4)=0\] for \(x\)

OpenStudy (anonymous):

since it is already factored you have very little work to do \[x-2=0\iff x=2\] so \(x=2\) is one answer also \[x+4=0\iff x=-4\] making your two answers \(x=2,x=-4\)

OpenStudy (anonymous):

Oh, so that's it? I was thrown off by it asking for zeros.. what does the multiplicity part mean?

OpenStudy (mathmale):

In your posted problem, (x-2)(x+4)=0, you'll notice that each of the factor (x-2) and (x+4) shows up just once. We say each has a "multiplicity" of one. Supposing you had (x-2)(x-2)(x+4)=0 and notice that the factor (x-2) shows up twice. We thus say that (x-2) has a multiplicity of two. Thinking about this a bit, I realize that the term "multiplicity" applies primarily to the roots (such as x=2, x=-4) more to the factors. So, if (x-2) shows up twice, the root 2 has a multiplicity of two.

OpenStudy (anonymous):

means the exponent of the factor, which is one in every case in your example

OpenStudy (anonymous):

\[(x-3)^4\] has a zero of \(3\) with multiplicity \(4\)

OpenStudy (anonymous):

ohhh ok. what about y=x^2(x+1)? how would I do that

OpenStudy (the_fizicx99):

y = x^2(x + 1) x^2 = 0 <-- ._. basic math it's zero >.> x + 1 = 0 -1 -1 x = -1